What is the pH change of a 0.300 M solution of citric acid (pKa=4.77) if citrate is added to a concentration of 0.145 M with no change in volume?
Answer #1

Given citric acid concentration = 0.3 M
pKa = 4.77 ==> Ka = 10-pKa = 10-4.77 = 0.00001698243
[H+] = sqrt(Ka x conc) = sqrt(0.00001698243 x 0.3) = 0.00000509473
pH = -log[H+] = -log[0.00000509473] = 5.29
pH = 5.29

before addition of citrate pH = 5.29

after addition of citrate

fente Hend 今uation ha,fo be used Os 2

after addition of citrate pH = 4.45
change in pH = 5.29 - 4.45 = 0.84

Final Answer change in pH = 0.84


Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")